OpenStudy (anonymous):

The permeter of right triangle ABC is 132 Find the length of the hypotenuse .... AB is 4x+17 and CB (bottom line) 5x+5 and AC = x ???

OpenStudy (bahrom7893):

Okay so: P = 132 = AB + AC + CB = 4x + 17 + 5x + 5 + x

OpenStudy (bahrom7893):

132 = 10x + 22

OpenStudy (bahrom7893):

132-22 = 10x

OpenStudy (bahrom7893):

110 = 10x; x = 11

OpenStudy (bahrom7893):

So: AB = 44+17 = 61 CB = 55+5 = 60

OpenStudy (bahrom7893):

AC = 11

OpenStudy (bahrom7893):

Hypothenuse is 61

OpenStudy (bahrom7893):

please click on become a fan if I helped! Thanks!

OpenStudy (anonymous):

thank you!!!!

OpenStudy (bahrom7893):

please click on Become a fan if I helped, thanks! =)